# Well-posedness of a generalized Stokes operator on domains with cylindrical ends via layer-potentials

### 11 May 2026 v1

@@ 10.2. Invertibility of $\frac{1}{2}+\boldsymbol{K}$
The proof of the invertibility of $\frac{1}{2}+\boldsymbol{K}$ (under suitable conditions on $V$ and $V_{0}$) requires the invertibility of its indicial operators. We first consider the case $\tau=0$. This case needs additional assumptions and arguments. Recall that we have $V(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)$ for $(x^{\prime},t)\in M^{\prime}\times(-\infty,0)$ and $t\ll 0$ and that $\color[rgb]{1,0,1}\widehat{V}(x^{\prime},t)=\color[rgb]{1,0,1}\widehat{V}(x^{\prime})$ is independent of $t$. The function $\color[rgb]{1,0,1}\widehat{V}_{0}$ on $M^{\prime}$ is defined similarly. The functions $\color[rgb]{1,0,1}\widehat{V}_{0}$ and $\color[rgb]{1,0,1}\widehat{V}$ correspond to the limit operators of the operators defined by $V_{0}$ and $V$.
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions (Assumption 8.3) and let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (10.2)). We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ on one side of its boundary $\Gamma^{\prime}:=\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Of course, in view of our conventions, the condition $\color[rgb]{1,0,1}\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
@@
The proof is very similar to that of Proposition 10.1, but using Proposition 9.10 instead of 9.7. Indeed, first, the hypotesis imply that the operator $\boldsymbol{\Xi}$ is Fredholm, and hence the layer potentials are defined. The result then follows from Proposition 9.10 applied to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ acting on $M^{\prime}\times\mathbb{R}$ and the fact that layer potentials commute with limit operators, Theorem 6.10. ∎
@@
Let us assume that $V,V_{0}\geq 0$ satisfy Assumption 8.3 (so $\color[rgb]{1,0,1}\widehat{V}\succ 0$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $M^{\prime}$)and let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (10.2)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 and let $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ be as in Theorem 8.14 (as before, see also Equation (10.2)). We furthermore assume that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator on $L^{2}(\Gamma;TM)$.
@@
The proof is similar to that of Corollary 10.3, but using Theorem 9.12 instead of Theorem 9.9. Indeed, the operator ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$ is the double layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\color[rgb]{1,0,1}\widehat{V},\color[rgb]{1,0,1}\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.12 are therefore satisfied by ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$, which is hence invertible. Because $\frac{1}{2}+\boldsymbol{K}$ is elliptic, Theorem 6.15 gives that $\frac{1}{2}+\boldsymbol{K}$ is Fredholm. ∎
@@
Let us assume that $M,V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $M^{\prime}$), that $\Omega$ is connected, and that $\Omega$ is the interior of $\overline{\Omega}$. We furthermore assume the following:
@@
    $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$,
@@
    $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and
@@
Step 2 (Single layer $U:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$). To prove that $\frac{1}{2}+\boldsymbol{K}$ is onto, it suffices to show that $\ker(\frac{1}{2}+\boldsymbol{K}^{*})=0$. To this end, let $\boldsymbol{h}\in L^{2}(\Gamma;TM)$ be such that $(\frac{1}{2}+\boldsymbol{K}^{*})\boldsymbol{h}=0$. The ellipticity of $\frac{1}{2}+\boldsymbol{K}^{*}$, the fact that $\Gamma$ has straight cylindrical ends, and the elliptic regularity property ([67, Proposition 6.5]) imply that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Let now the single layer potential $\mathcal{S}_{\rm{ST}}\boldsymbol{h}$. In view of Proposition 8.9 we obtain that the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy
@@
Step 4 (Study of $U$ on $\Omega_{+}$). Let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$, as in the previous two steps. Recall that $\boldsymbol{\Xi}U=0$ in $M\smallsetminus\Gamma$. Theorem 8.16 gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma:=\partial\Omega$ of $\Omega$ (interior and exterior traces). The fact that $\boldsymbol{u}_{+}=0$ at $\Gamma$ allows us to use Corollary 2.14 that gives then that $\boldsymbol{u}=0$ in $\Omega$. Because $\Omega$ is connected and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, we also have $V_{0}\succ 0$ on $\Omega$. Therefore, the same corollary implies that $p$ is constant on $\Omega$. Because $V_{0}$ does not vanish identically on $\Omega$ (as we have just noticed) we even obtain that $p=0$ in $\Omega$ and $p_{+}=0$. Recalling that we have already proved that $\boldsymbol{u}=0$ in $\Omega_{-}$, we see that $\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$ and hence $\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$.
@@
The fact that $(\frac{1}{2}+\boldsymbol{K})^{-1}\in\Psi^{0}(\Gamma;TM)$ follows from Theorem 5.4.11 in [63] (see also Theorem 7.4 in [67]). This completes the proof. ∎
@@ 11.1. The homogeneous Stokes system
The invertibility of each of the operators $\frac{1}{2}+\boldsymbol{K}$ and $\boldsymbol{S}$ gives the following well-posedness result for the modified Stokes system with Dirichlet boundary condition, problem (1.4). This result is known for the usual Stokes operator (i.e. when $V=0$ and $V_{0}=0$, see [100, Proposition 10.5.1, Theorem 10.6.2] in the case of a bounded Lipschitz domain in $\mathbb{R}^{n}$, $n\geq 2$, and [39, Theorem 5.1] in the case of a $C^{1}$ domain on a compact manifold, see also [63] for further related applications). We consider the homogeneous Dirichlet problem
@@
Let $M$, $V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular, $M$ is has straight cylindrical ends) and let $\Omega\subset M$ be a compatible smooth domain (so $\Omega$ also has straight cylindrical ends, see Equation (1.2)). We assume that $V_{0}\succ 0$ on $\Omega$, that $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and that $\Omega$ is on one side of its boundary $\Gamma:=\partial{\Omega}$. We further assume one of the following conditions:
@@
    $V\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ or
@@
    $V_{0}\succ 0$ on $\Omega_{-}$ and $\color[rgb]{1,0,1}\widehat{V}_{0}\succ 0$ on $\Omega_{-}^{\prime}$.
@@
Our assumption (a) implies that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8. Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$.
If, on the other hand, condition (b) is satisfied, then the operator $\boldsymbol{S}$ is invertible on $L^{2}(\Gamma;TM)$ and $\boldsymbol{S}^{-1}\in\Psi^{1}(\Gamma;TM)$, by Theorem 10.4. Moreover, the elliptic regularity of the operator $\boldsymbol{S}:=\mathcal{A}_{0}$ (Theorem 8.16) implies that $\boldsymbol{h}:=\boldsymbol{S}^{-1}\boldsymbol{f}\in H^{m-1/2}(\Gamma;TM)$. Theorem 6.6 then gives
@@
Next we prove the existence part of Theorem 1.1. Given the potential $V$ satisfying the assumptions of this theorem, we can modify it on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$. Then our assumptions imply that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8. Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U_{1}:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$. This completes the proof. ∎
@@
We notice that the solution $U$ of the above theorem can be represented either by a single or as a double layer potential, depending on how we modify $V$ and $V_{0}$ on $\Omega_{-}$. Indeed, modifying $V$ on $\Omega_{-}$ such that $\color[rgb]{1,0,1}\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$, then the assumptions of Theorem 10.8 are satisfied and the solution $U$ can be represented as a double layer potential
@@
The constant $\mathcal{C}_{m}$ in Theorem 1.1 is independent of the given datum $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Indeed, it depends only on the norms of the operators $\mathcal{D}_{\rm{ST}}$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}$ and can be chosen the product of them.

### 28 May 2026 v2

@@ 10.2. Invertibility of $\frac{1}{2}+\boldsymbol{K}$
The proof of the invertibility of $\frac{1}{2}+\boldsymbol{K}$ (under suitable conditions on $V$ and $V_{0}$) requires the invertibility of its indicial operators. We first consider the case $\tau=0$. This case needs additional assumptions and arguments. Recall that we have $V(x^{\prime},t)=\widehat{V}(x^{\prime},t)$ for $(x^{\prime},t)\in M^{\prime}\times(-\infty,0)$ and $t\ll 0$ and that $\widehat{V}(x^{\prime},t)=\widehat{V}(x^{\prime})$ is independent of $t$. The function $\widehat{V}_{0}$ on $M^{\prime}$ is defined similarly. The functions $\widehat{V}_{0}$ and $\widehat{V}$ correspond to the limit operators of the operators defined by $V_{0}$ and $V$.
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions (Assumption 8.3) and let $\widehat{\boldsymbol{K}}(0)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$ given in Theorem 8.14 (see also Equation (10.2)). We furthermore assume that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ on one side of its boundary $\Gamma^{\prime}:=\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\widehat{\boldsymbol{K}}(0)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(0)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
Of course, in view of our conventions, the condition $\widehat{V}_{0}\not\equiv 0$ on $\Omega^{\prime}$ implies the condition $V_{0}\not\equiv 0$ on $\Omega$ (since $V$ and $V_{0}$ are translation invariant at infinity).
@@
The proof is very similar to that of Proposition 10.1, but using Proposition 9.10 instead of 9.7. Indeed, first, the hypotesis imply that the operator $\boldsymbol{\Xi}$ is Fredholm, and hence the layer potentials are defined. The result then follows from Proposition 9.10 applied to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi})$ acting on $M^{\prime}\times\mathbb{R}$ and the fact that layer potentials commute with limit operators, Theorem 6.10. The last point follows from Theorem 6.17. ∎
@@
Let us assume that $V,V_{0}\geq 0$ satisfy Assumption 8.3 (so $\widehat{V}\succ 0$ and $\widehat{V}_{0}\succ 0$ on $M^{\prime}$)and let $\widehat{\boldsymbol{K}}(\tau)\in\Psi^{0}(\Gamma^{\prime};TM)$ be the indicial operator of the operator $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ of Theorem 8.14 (see also Equation (10.2)). If $\tau\neq 0$, then $\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)$ is invertible on $L^{2}(\Gamma^{\prime};TM)$ and $\big[\frac{1}{2}+\widehat{\boldsymbol{K}}(\tau)\big]^{-1}\in\Psi^{0}(\Gamma^{\prime};TM)$.
@@
Let us assume that $M,V$, and $V_{0}$ satisfy the non-vanishing assumptions 8.3 and let $\boldsymbol{K}\in\Psi^{0}(\Gamma;TM)$ be as in Theorem 8.14 (as before, see also Equation (10.2)). We furthermore assume that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and that $\Omega^{\prime}$ is the interior of $\overline{\Omega^{\prime}}$. Then $\frac{1}{2}+\boldsymbol{K}$ is a Fredholm operator on $L^{2}(\Gamma;TM)$ and its pseudo-inverse $(\frac{1}{2}+\boldsymbol{K})^{(-1)}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$.
@@
The proof is similar to that of Corollary 10.3, but using Theorem 9.12 instead of Theorem 9.9. Indeed, the operator ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$ is the double layer potential operator associated to ${\mathcal{R}_{\infty}}(\boldsymbol{\Xi}_{V,V_{0}})=\boldsymbol{\Xi}_{\widehat{V},\widehat{V}_{0}}$, by Theorem 6.10. The assumptions of the corollary give that the assumptions of Theorem 9.12 are therefore satisfied by ${\mathcal{R}_{\infty}}(\frac{1}{2}+\boldsymbol{K})$, which is hence invertible. Because $\frac{1}{2}+\boldsymbol{K}$ is elliptic, Theorem 6.15 gives that $\frac{1}{2}+\boldsymbol{K}$ is Fredholm. The last point follows from Theorem 6.17. ∎
@@
Let us assume that $M,V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular $\widehat{V}\succ 0$ on $M^{\prime}$), that $\Omega$ is connected, and that $\Omega$ is the interior of $\overline{\Omega}$. We furthermore assume the following:
@@
    $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$,
@@
    $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$, and
@@
Step 2 (Single layer $U:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$). To prove that $\frac{1}{2}+\boldsymbol{K}$ is onto, it suffices to show that $\ker(\frac{1}{2}+\boldsymbol{K}^{*})=0$. To this end, let $\boldsymbol{h}\in L^{2}(\Gamma;TM)$ be such that $(\frac{1}{2}+\boldsymbol{K}^{*})\boldsymbol{h}=0$. The ellipticity of $\frac{1}{2}+\boldsymbol{K}^{*}$, the fact that $\Gamma$ has straight cylindrical ends, and the elliptic regularity property (see Theorem 6.16 or [83, Theorem 15.3.30] or [87, Proposition 6.5]) imply that $\boldsymbol{h}\in H^{s}(\Gamma;TM)$ for all $s\in\mathbb{R}$. Let now the single layer potential $\mathcal{S}_{\rm{ST}}\boldsymbol{h}$. In view of Proposition 8.9 we obtain that the restrictions of $U$ to $\Omega_{+}:=\Omega$ and to $\Omega_{-}$ satisfy
@@
Step 4 (Study of $U$ on $\Omega_{+}$). Let $U=(\,\boldsymbol{u}\ \ p\,)^{\top}:=\mathcal{S}_{\rm{ST}}(\boldsymbol{h})$, as in the previous two steps. Recall that $\boldsymbol{\Xi}U=0$ in $M\smallsetminus\Gamma$. Theorem 8.16 gives the “no-jump relation” $\boldsymbol{u}_{+}=\boldsymbol{u}_{-}=0$ at the boundary $\Gamma:=\partial\Omega$ of $\Omega$ (interior and exterior traces). The fact that $\boldsymbol{u}_{+}=0$ at $\Gamma$ allows us to use Corollary 2.14 that gives then that $\boldsymbol{u}=0$ in $\Omega$. Because $\Omega$ is connected and $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, we also have $V_{0}\succ 0$ on $\Omega$. Therefore, the same corollary implies that $p$ is constant on $\Omega$. Because $V_{0}$ does not vanish identically on $\Omega$ (as we have just noticed) we even obtain that $p=0$ in $\Omega$ and $p_{+}=0$. Recalling that we have already proved that $\boldsymbol{u}=0$ in $\Omega_{-}$, we see that $\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$ and hence $\boldsymbol{D}_{\boldsymbol{\nu}}\boldsymbol{u}=0$ in $M\smallsetminus\Gamma$.
@@
The fact that $(\frac{1}{2}+\boldsymbol{K})^{-1}\in\Psi^{0}(\Gamma;TM)$ follows from Theorem 6.17 (see Theorem 7.4 in [87] or Theorem 15.4.11 in [83]). This completes the proof. ∎
@@ 11.1. The homogeneous Stokes system
The invertibility of each of the operators $\frac{1}{2}+\boldsymbol{K}$ and $\boldsymbol{S}$ gives the following well-posedness result for the modified Stokes system with Dirichlet boundary condition, problem (1.5) (see Equation (1.1) for a more explicit form of this problem). This result is known for the usual Stokes operator (i.e. when $V=0$ and $V_{0}=0$, see [122, Proposition 10.5.1, Theorem 10.6.2] in the case of a bounded Lipschitz domain in $\mathbb{R}^{n}$, $n\geq 2$, and [49, Theorem 5.1] in the case of a $C^{1}$ domain on a compact manifold, see also [83] for further related applications). We consider the homogeneous Dirichlet problem
@@
Let $M$, $V$, and $V_{0}$ satisfy Assumption 8.3 (so, in particular, $M$ is has straight cylindrical ends) and let $\Omega\subset M$ be a compatible smooth domain (so $\Omega$ also has straight cylindrical ends, see Equation (1.3)). We assume that $V_{0}\succ 0$ on $\Omega$, that $\widehat{V}_{0}\succ 0$ on $\Omega^{\prime}$, and that $\Omega$ is on one side of its boundary $\Gamma:=\partial{\Omega}$. We further assume one of the following conditions:
@@
    $V\succ 0$ on $\Omega_{-}$ and $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ or
@@
    $V_{0}\succ 0$ on $\Omega_{-}$ and $\widehat{V}_{0}\succ 0$ on $\Omega_{-}^{\prime}$.
@@
Our assumption (a) implies that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi_{\operatorname{ess}}^{0}(\Gamma;TM)$, by Theorem 10.8. The mapping properties of the $\operatorname{ess}$-calculus” then gives that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$.
If, on the other hand, condition (b) is satisfied, then the operator $\boldsymbol{S}$ is invertible on $L^{2}(\Gamma;TM)$ and $\boldsymbol{S}^{-1}\in\Psi_{\operatorname{ess}}^{1}(\Gamma;TM)$, by Theorem 10.4. The mapping properties of the $\operatorname{ess}$-calculus” then imply that $\boldsymbol{h}:=\boldsymbol{S}^{-1}\boldsymbol{f}\in H^{m-1/2}(\Gamma;TM)$. Theorem 6.6 then gives
@@
Next we prove the existence part of Theorem 1.1. Given the potential $V$ satisfying the assumptions of this theorem, we can modify it on $\Omega_{-}$ such that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$. Then our assumptions imply that the operator $\frac{1}{2}+\boldsymbol{K}$ is invertible on $L^{2}(\Gamma;TM)$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}\in\Psi^{0}(\Gamma;TM)$, by Theorem 10.8. Moreover, the elliptic regularity of the operator $\frac{1}{2}+\boldsymbol{K}$ (Theorem 8.14) implies that $\boldsymbol{h}:=(\frac{1}{2}+\boldsymbol{K})^{-1}\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Then $U_{1}:=\mathcal{D}_{\rm{ST}}(\boldsymbol{h})$ satisfies the required properties, including its boundedness with respect to the given datum $\boldsymbol{f}$, with a constant $C_{m}$ given by $\|\mathcal{D}_{\rm{ST}}(\frac{1}{2}+\boldsymbol{K})^{-1}\|$. This completes the proof. ∎
@@
We notice that the solution $U$ of the above theorem can be represented either by a single or as a double layer potential, depending on how we modify $V$ and $V_{0}$ on $\Omega_{-}$. Indeed, modifying $V$ on $\Omega_{-}$ such that $\widehat{V}\succ 0$ on $\Omega_{-}^{\prime}$ and $V\succ 0$ on $\Omega_{-}$, then the assumptions of Theorem 10.8 are satisfied and the solution $U$ can be represented as a double layer potential
@@
The constant $\mathcal{C}_{m}$ in Theorem 1.1 is independent of the given datum $\boldsymbol{f}\in H^{m+1/2}(\Gamma;TM)$. Indeed, it depends only on the norms of the operators $\mathcal{D}_{\rm{ST}}$ and $\big[\frac{1}{2}+\boldsymbol{K}\big]^{-1}$ and can be chosen the product of them. See also [3] and [4] for other results related to the Stokes operator on unbounded domains.